OpenStudy (anonymous):

You roll two dice. What is the probability of the event that you roll a total of three or eleven on the two dice?

OpenStudy (anonymous):

you would get a probability of 1/9 the only combinations that will get you 3 or 11 are (1,2) (2,1) (5,6) and (6,5)

OpenStudy (anonymous):

= pr(3) + pr(11) pr(3) = pr(2)*pr(1) + pr(1)*pr(2) pr(11) = pr(5)*pr(6) + pr(6)*pr(5) = (1/6)^2 + (1/6)^2 + (1/6)^2 + (1/6)^2 = 4*(1/6)^2 = 4/36 = 1/9